OpenStudy (anonymous):

Evaluate the integral of sqrt(x^2-121) divided by x^3 with respect to x.

OpenStudy (anonymous):

with respect to x?

OpenStudy (anonymous):

\[\int\limits_{}^{}(\sqrt{x^2 - 121})/x^3 dx\]

OpenStudy (anonymous):

However you word dx...

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{x^2-121}{x^3}dx=\int\limits_{}^{}(x^2-121)(x^{-3})\]

OpenStudy (anonymous):

dx

OpenStudy (anonymous):

x^2-121 is under a square root sign

OpenStudy (anonymous):

trig sub then

OpenStudy (anonymous):

let x=asec(x)

OpenStudy (anonymous):

so x=11sec(theta)

OpenStudy (anonymous):

think you can handle it from there?

OpenStudy (anonymous):

That's what I did, and I got: \[(1/22)\sec^{-1}(x/11)-\sqrt{x^2-121}/2x^2 +C \] but my online homework isn't taking it :(

OpenStudy (anonymous):

I don't know if the way I typed it is incorrect, but I just want to check if someone else gets the same answer.

OpenStudy (anonymous):

what did you get when you subbed everything in

OpenStudy (anonymous):

and cancelled everything

OpenStudy (anonymous):

\[1/11\int\limits_{}^{}\tan^{2} \theta/\sec^{2} \theta d{\theta}\]

OpenStudy (anonymous):

correct so far

OpenStudy (anonymous):

Then: \[1/11\int\limits_{}^{}\sin^{2}{\theta}/\cos^{2}{\theta} * \cos^{2}{\theta} d{\theta}\]

OpenStudy (anonymous):

(1/sec^2(theta) = cos^2(theta) )

OpenStudy (anonymous):

Cancel cosines and get 1/11 integrand sin^2(theta) d(theta)

OpenStudy (anonymous):

alright so far so good

OpenStudy (anonymous):

then I made sin^2 to 1-cos^2

OpenStudy (anonymous):

oh and 1/2 on the outside

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

So:\[(1/11)*(1/2)\int\limits_{}^{}(1-\cos^{2}{\theta})d{\theta}\] \[(1/22)\int\limits_{}^{}(1-\cos^{2}{\theta})d{\theta}\] \[(1/22)[{\theta}-(1/2)\sin(2{\theta})] +C\]

OpenStudy (anonymous):

Then to convert back to x's, used a right triangle.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

in your above answer at the top where did you get x^2

OpenStudy (anonymous):

\[{\theta}=\sec^{-1}(x/11)\] \[\sin{2}{\theta}=2\sin{\theta}\cos{\theta}=(\sqrt{x^2-121}/x)*(11/x)\]

OpenStudy (anonymous):

Then times sin2theta by 1/2 Then times both by 1/22 :P

OpenStudy (anonymous):

so you got sin(2theta) = 2sinthetacostheta \[\frac{2}{1}\frac{\sqrt{x^2-121}}{x}*\frac{11}{x}=\frac{22}{1}\frac{\sqrt{x^2-121}}{x^2}\]

OpenStudy (anonymous):

Times 1/2, cuz originall it was (1/2)sin^2(theta)

OpenStudy (anonymous):

so the 2's cancel= (1/2)*2*sin(theta)cos(theta)

OpenStudy (anonymous):

correct hmm

OpenStudy (anonymous):

I tried entering it again... I can only submit it twice, but it worked the 2nd time :P I guess it didn't like (sec(x/11)^-1 at first so I used their built-in sec^-1. Ugh so frustrating, but at least I know I'm doing the steps correctly. Thanks for checking.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

if it didn't work i would have argued to teacher taht the answer was correct

OpenStudy (anonymous):

because that answer is correct

OpenStudy (anonymous):

Do you know anything about cylindrical shells? I have an unanswered question I posted earlier :)

OpenStudy (anonymous):

cylindrical shells?

OpenStudy (anonymous):

volume by c.s.